Two sides of a triangle measure 15 inches and 18 inches, respectively. Which of these is NOT a possible length for the third side of the triangle?
Answers available in ...
10Take this time to do your
best on this question.
A)
4 inches
B)
18 inches
C)
32 inches
D)
36 inches

Found 2 solutions by katealdridge, rwm:
Answer by katealdridge(100)

If you want to find the possibilities for the third side, add the numbers together:15+18=33. Then subtract the numbers 18-15=3. The third side must be between these two numbers but not including these numbers. Therefore the only number that does not fit into this range is 36. Remember that the third side could NOT be 3 or 33.

Answer by rwm(914)

The sum of any two sides must be greater than the third side.
15+18=33
so the third side must be less than 33 which excludes d 36